OpenStudy (anonymous):

given the linear equation y=1/2x-5 find the y coordinates of the points (-8), (-2)and (4). show work

OpenStudy (anonymous):

1/-21,1/-9,1/3

OpenStudy (amistre64):

just gotta plug in the numbers to get the results...

OpenStudy (anonymous):

im sorry i still dont understand

OpenStudy (anonymous):

Just plug in the values for x y= 1/2*(-8)-5 = 1/-21 y= 1/2*(-2)-5 = 1/-9 y= 1/2*(4)-5 = 1/3 the numbers that you plug in are in brackets Hope this helps

OpenStudy (anonymous):

\[y=\frac{1}{2*(-8)-5}\] \[y=\frac{1}{2*(-2)-5}\] \[y=\frac{1}{2*(-2)-5}\]

OpenStudy (anonymous):

yeah latex didn't work out for me more practice needed please ignore last entry :D